If the Present value is (Round your response to the nearest penny) Definition Present Value Todays value of a payment to be received in the future when the interest rate is i Print Done
Answer #1

Solution: 2755

Working:

PV = FV1/(1+i) + FV2 /(1+i)^2 + FV3 /(1+i)^3

= 1,100/ 1.15 + 1,220 / 1.3225 + 1,333 / 1.520875

= 956.52 + 922.50 + 876.47

= 2755.49